CLEAR Weekly Operation Meeting

Europe/Zurich
CERN

CERN

Videoconference Rooms
CLEAR Weekly Operation Meeting
Zoom Meeting ID
63110816470
Host
Roberto Corsini
Useful links
Join via phone
Zoom URL

MINUTES

Actions from previous week:

  • Check the availability of a beam pipe to allow the installation of only two CLIC cavity BPM --> Stefano has had no contact, the usual confusion over this project. 
  • Further look to solve the noise in the plasma lens pump --> this is an ongoing task, but no obvious noise was heard when we accessed.
  • Devise a way of doing laser alignment on the in-air test stand. --> I believe Wilfrid ordered the laser, but the system has still not been installed.
  • Understand where's the end of the trigger cables in the klystron gallery. --> Pretty much done, but Antonio suggests that its worth keeping on the actions.
  • Look at moving one of the unused CTF3 front ends to the barrack to aid data collection. --> Davide has received no news and will follow it up.
  • Reconnection of the thermalised water pipes on the X-band ACS --> Connected, although there was a control issue.  The person responsible previously has retired.  Antonio does not know how to open the valves, although he thinks he now knows where the controls are.  For this week's operation they do not think that they will need to use the pumps, unless the temperature rises above a certain level.
  • Re-installation of safety valve on gas pipe for P.L.    --> ongoing, long term
There are minutes attached to this event. Show them.